PREX/CREX May 7, 2019 10:00am

From PREX Wiki
Jump to: navigation, search

Back to Main Page >> PREX/CREX_Teleconferences

previous meeting << >> following meeting

Logistic information

 BlueJeans calling instructions:
 Toll-Free Number (U.S.&  Canada):  888-240-2560
 International toll number:         408-740-7256
 Bluejeans CODE:                  628 981 532
 Bluejeans link: https://bluejeans.com/628981532
 Offline room:   CEBAF Center F324-325

Agenda

  1. Installation progress: Ciprian docDB363
    • CG: KP and SR is evaluating the survey results for the vacuum enclosures and their impact on our FOM.
  2. Compton polarimeter
    • DG: we are still trying to understand the cavity birefringence. The amp is in shipping and receiving
  3. Moller polarimeter
    • EK: The commissioning results are consistent with the simulation results. We have a high stats for the PREX tune and we will compare with G3 (the differences are expected in the Levchuk corrections).
    • DJ: EK and PS have come up with a strategy to match the simulation and data using the asymmetry (including the Levchuk effect) -- effectively tuning Q1 and showing that Q2 and Q4 and insensitive to 10% detunes. This will be crucial in order to obtain a smaller systematic.
  4. BPM studies
    • CP: working on elog. Produced a type a noise that is synced with 60Hz noise. Can see double peaks in normal mode and no peak in the fast feedback mode — reproduces the noise we see with beam.
  5. Integrating DAQ updates
    • BM: we switched the parity adaq1 and LHRS daq to adaq3.
    • BM: ChG, YT, DA working on the HRS daq. We are reviving the crates for the GEMs.
    • PK: on the integrating DAQ CC and I realized we need a timing board for the VQWKs to support the tune mode scans. We are testing and getting that working.
    • CC: finished testing all the VME boards. We have 5 VQWKs and 2 FLEXIOs spares.
  6. Parity analyzer:
    • PK: we need to define the size of the blinding box. In the past we (Qweak) used 10x the projected error bar.
  7. LQ1 viton seals:
    • AZ: latest simulations on haplog.
  8. Injector studies:
    • CP: we are making progress on pickup issues. It’s at the ppb level in the crate. on the injector table we saw ppm pickup. We can reduce the pickup with a different configuration.
    • PK: we should run a battery from the laser table to upstairs at least for a short test. KK: will get involved in this as well and we’ll discuss soon.
  9. Detectors:
    • SR: Chandan assembled the big frames. We are going to leave the rest to the techs. Jack said today might be the earliest to put in the frames. ChG: today we will move 1 SBS GEM to the hall. He will crane the detectors in the afternoon at the earliest.
  10. Discussion for next week:
    • EK: will present Moller polarimeter update in 2 weeks.

Present

In F324

KK, Adam, Tao, Caryn, Ye T, Wouter D, Devi A

Online

Sanghwa P, Paul K, Guido U, Ciprian G, Seamus R, Dave G, Eric K, Victoria O, Ryan R, Chandan G, Cameron C, Jay B, Paul S, Amali P, Juan Carlos C, Robert R, Jinlong Z, Silviu D, Don Jones, Nilanga L, Rich H, Bob M, Sakib R, Siyu J

Cannot make it

Kent P,